# Identification of malfunctioning quantum devices

### v2

@@ I Introduction
Recent advancements in quantum technologies, such as quantum computing devices Nigg et al. (2014); Taminiau et al. (2014); Córcoles et al. (2015); Kelly et al. (2015); Náfrádi et al. (2016); Cramer et al. (2016), quantum communication Yin et al. (2017); Ren et al. (2017); Liao et al. (2017), and quantum sensors Reiter et al. (2017), lend credence to the notion that one day soon such devices will be readily available and, hopefully, part of an interconnected quantum network Kimble (2008). In turn, the existence of such quantum networks gives rise to new technical challenges, such as the correct identification of possible malfunctions. In a vast network of quantum devices—be they sources that produce quantum states, quantum channels that transmit information, or the vast array of gates in quantum computers—it is imperative that we are able to find efficient ways to identify faulty components.
@@
The identification of rare events that differ significantly from the majority of all other observations is known as anomaly detection and is a fundamental primitive in classical data analysis and signal processing Tartakovsky et al. (2015) with a wide range of applications from the identification of denial of service attacks Thottan and Ji (2003); Siris and Papagalou (2004); Tartakovsky et al. (2013), to the identification of fraudulent financial activity Ahmed et al. (2016) (for a survey of anomaly detection see Chandola et al. (2009)). Two important algorithms for anomaly detection, namely kernel principal component analysis and support vector machines, have been shown to be efficiently applicable in detecting anomalies in quantum data Liu and Rebentrost (2018). Here we consider a more direct application of anomaly detection where the anomaly is not restricted to classical data but in the most general quantum device. Specifically, given $N$ identical devices programmed to perform a particular task, with one of the devices developing a known malfunction, our goal is to optimally identify the malfunctioning device whilst only allowed to query the network once (see Fig. 1). This task that we term Error Position Identification, or EPI for short, is a key ingredient in pulse-position modulated quantum communication Dolinar et al. (2006); Dalla Pozza and Pierobon (2015), quantum illumination Lloyd (2008), quantum reading Pirandola (2011), and target detection Zhuang et al. (2017).
@@
We note that instances of EPI have appeared elsewhere in the literature under the name of channel position finding Zhuang and Pirandola (2020a, b); Harney and Pirandola (2021); Pereira et al. (2021) where the task is to identify the position of a target bosonic channel, with a given reflectivity, among $m$ background bosonic channels of a different reflectivity. The case of identifying the position of a singular unitary channel was also considered recently in Hillery et al. (2023). Here, we provide a more comprehensive analysis of EPI by considering more general devices—including states, unitary gates, and quantum channels—using a variety of different strategies employing probes in separable, entangled, as well as ancilla assisted strategies. Unlike Zhuang and Pirandola (2020a, b); Harney and Pirandola (2021); Pereira et al. (2021) we consider channels acting on two-dimensional systems, and consider both rank-one and rank-two Pauli noise channels, depolarizing, and amplitude damping channels. For the case of unitary channels we show that there exists instances for which the position of the unitary channel can be identified with certainty, a result that was missed in Hillery et al. (2023). Note that EPI is fundamentally different from the anomaly detection scenario considered in Liu and Rebentrost (2018) which focused in classification of states using quantum machine learning techniques.
@@
We first address the setting when the devices are quantum sources known to produce a specific pure state in $\mathcal{H}_{d}$ with the faulty source producing a different anomalous pure quantum state (the multi-anomaly case has recently been studied in Llorens et al. (2023)). We show that the maximum probability of success is achieved by a global measurement strategy given by the so-called square root measurement, which yields a success probability that converges to a non-zero constant as $N$ increases. We next show that when the devices are known to perform a specific unitary gate, with the anomalous device performing a different unitary, a probe state with entanglement among the $N$ input states—that can be effectively taken to be qubits—achieves the maximal probability of success without the need of using ancillary systems. In fact, we show that for certain non-trivial errors the optimal probability of success can even reach unity. Whilst the use of entangled probe states does yield an improvement over strategies employing separable probe states, this improvement diminishes with the size of the network.
@@
The task of successfully identifying the position of a faulty device is equivalent to optimally discriminating among $N$ quantum states, in the case of source EPI, or $N$ quantum channels. For the remainder of this work we will be concerned with devices that either prepare systems in a known pure state (sources), or perform a known quantum operation (channels). Without loss of generality a faultless quantum source prepares systems in the state $\left|0\right\rangle$ whereas the faulty source—equally likely to be located at any of the $N$ positions—prepares the $k^{\mathrm{th}}$ system in some state $\left|\phi\right\rangle$. Similarly, a faultless quantum operation performs the unitary operation $V$ whereas the anomalous operation performs a completely-positive, trace-preserving (CPTP) map, $\mathcal{F}^{(k)}$, on the $k^{\mathrm{th}}$ quantum system  [^1] In the language of Liu and Rebentrost (2018) EPI is an instance of point anomaly and, moreover, can be cast in the framework of supervised learning as we assume that the description of both normal and anomalous devices is known. The probability of successfully identifying the position of the faulty device is given by
@@
where $q(l|k)$ is the conditional probability given by $\mathrm{Tr}\left[M_{l}\,\left|\phi_{k}\right\rangle\!\!\left\langle\phi_{k}\right|\right]$ in the case of sources, and $\mathrm{Tr}\left[M_{l}\,\mathcal{E}^{(k)}[\left|\psi\right\rangle\!\!\left\langle\psi\right|]\right]$ in the case of channels, with $\{M_{l}\geq 0\,|\sum_{l}M_{l}={\mbox{$1{\bf l}$}}\}:=\mathcal{M}$ forming a general positive operator-valued measure (POVM). Observe that for source EPI, the optimization over all measurement strategies of Eq. 1 is a semi-definite program (SDP). For channel EPI, however, Eq. 1 needs to be optimized both over the input state as well as the measurement. Luckily enough, the formalism of quantum testers Chiribella et al. (2008) allows one to linearize the problem and formulate it as an SDP—as long as the optimization is carried over strategies that make use of idler or ancillary systems as shown in Fig. 1.
By and large analytical solutions to either state or channel discrimination problems are very difficult and known only for two Holevo (1973); Helstrom (1976) or three Ha and Kwon (2013) pure or mixed states or between states possessing a certain symmetry Yuen et al. (1975); Ban et al. (1997); Eldar and Forney (2001); Eldar et al. (2004); Barnett (2001); Andersson et al. (2002); Sentís et al. (2016, 2017, 2018) (see also Barnett and Croke (2009); Bae and Kwek (2015)). For channel discrimination analytic results are known for distinguishing among a finite number of unitaries in a single or finite number of runs Acín (2001); Duan et al. (2007, 2009), or between two arbitrary CPTP maps Sacchi (2005a, b); Piani and Watrous (2009). We stress that here, unlike most channel discrimination instances to date, the identity of the channel (unitary or otherwise) is known, and what one is looking for is the position at which the channel is acting.
@@ III.1 Source EPI
Notice that the set of states also enjoys cyclic symmetry, which allows to work out the solution explicitly Ban et al. (1997); Eldar and Forney (2001); Eldar et al. (2004); Dalla Pozza and Pierobon (2015); Llorens et al. (2023).
Let $\{\left|m_{j}\right\rangle\}_{j=1}^{N}$ be an orthonormal basis for the $N$-dimensional space spanned by $\{\left|\psi_{k}\right\rangle\}_{k=1}^{N}$. Then the probability of success is given by Sentís et al. (2016)
@@
where $S$ is the unique, self-adjoint square root of $G$. The corresponding measurement is simply given by $\{\left|m_{k}\right\rangle=\frac{(VS)^{-1}}{\sqrt{N}}\left|\psi_{k}\right\rangle\}$ (see Sentís et al. (2016)).
@@
where $b=\left\langle\psi_{k}\lvert\psi_{l}\right\rangle$ for $k\neq l$ and $\left|\bm{1}\right\rangle=\frac{1}{\sqrt{N}}(1,\ldots,1)^{T}$. The matrix $G$ and has two distinct eigenvalues, $\lambda_{1}=\frac{1+(N-1)b}{N}$, and $\lambda_{2}=\frac{1-b}{N}$ where the latter is $(N-1)$-fold degenerate. Note that $G$ is circulant Horn and Johnson (2012) as is any function of $G$, in particular $S=\sqrt{G}$ whose diagonal entries are all equal to
@@
The condition that $S_{kk}=S_{ll},\,\forall k\neq l$ is necessary and sufficient to show that Eq. (3) is maximized by $B=S$ Dalla Pozza and Pierobon (2015); Sentís et al. (2016), i.e., $P_{S}^{*}=NS_{kk}^{2}=(\mathrm{Tr}{S})^{2}/N$, which reads
@@
The measurement that achieves this optimal value is the so-called square-root measurement Ban et al. (1997); Eldar and Forney (2001); Eldar et al. (2004); Dalla Pozza and Pierobon (2015). Notice that $P_{S}^{*}=1$ if and only if $b=0$, i.e., $\left|\phi\right\rangle=\left|1\right\rangle$. For large strings of states, the success probability is finite and reads
@@ III.2 Unitary EPI
where $U=V^{\dagger}W$, and $\left|\psi\right\rangle\in\mathcal{H}_{p}\otimes\mathcal{H}_{a}$ is a probe state which in principle can contain ancillary systems. However, we note that the use of ancillas here is redundant since as for any probe-plus-ancilla state there exists an $N$ probe state that gives the same Gram matrix (see Appendix A). Observe that if $\left|\psi\right\rangle=\left|\gamma\right\rangle^{\otimes N}$ with $U\left|\gamma\right\rangle=\left|\phi\right\rangle$, we recover the source scenario discussed above with success probability given by Eq. (7). The question is whether choosing a more suitable initial state, perhaps involving entanglement, improves the probability of success.
Using the symmetry of the problem, for any given input state $\rho\in\mathcal{B}(\mathcal{H}^{\otimes N})$ and optimal POVM $\{M_{\sigma}\}$ there exists a permutationally invariant state $\tilde{\rho}\in\mathcal{B}(\mathcal{H}^{\otimes N})$ and permutationally covariant POVM $\{\tilde{M}_{\sigma}\}$ that achieves the same probability of success (see Appendix B). On the other hand, as the probability of success is a convex function the optimal input state can always be chosen to be pure. A seemingly natural, though unproven, assumption is to restrict the optimization over states to those that are both pure and permutationally invariant. For $N>2$ this assumption amounts to searching over pure states that belong to the totally symmetric subspace of $\mathcal{H}^{\otimes N}$ (see Eq. 11). We have numerically verified that this is indeed the case for small values of $N$.
As $U$ is known we may write any permutationally invariant probe state with respect to the eigenbasis of $U$, with all amplitudes taken to be real and positive without loss of generality. Furthermore, an optimal probe state only involves those eigenstates of $U$ whose eigenvalues have the largest distance in the complex plane D’Ariano et al. (2001); Nakahira (2023) as these yield the smallest overlap. We write these as $\left|0\right\rangle$ and $\left|1\right\rangle$, respectively. The action of the unitary can be taken to be $U\left|0\right\rangle=\left|0\right\rangle$ and $U\left|1\right\rangle=e^{i\phi}\left|1\right\rangle$ without loss of generality, where $|\phi|$ is the largest phase distance among the eigenvalues of the unitary. To keep track of the phase we denote the unitary by $U(\phi)$.
An orthonormal basis of these permutationally symmetric states is given by the well known Dicke states Dicke (1954),
@@
We would like to point out that part of our results here are reproduced in a recent publication Hillery et al. (2023). Notwithstanding, this work assumes symmetric probe states of qubits ab initio and fails to notice that this probability can be made equal to one for non-trivial values of the angle $\phi$.
@@ III.3 Channel EPI
For rank-three Pauli channels, any strategy involving $N$ probes prepared in a product state and projective measurements is sub-optimal. This is because one cannot perfectly discriminate between the Pauli matrices using a single qubit. Indeed, the best one can hope for using such a strategy is (see Appendix C)
@@ IV Local Measurement Strategies for source EPI
In this section, we investigate four local measurement strategies and gauge their performance in determining the position of a malfunctioning quantum source, i.e. detecting faulty prepared states. For ease of exposition, we will describe each of the local measurement strategies and state their performance only. For a more detailed treatment of each strategy, we direct the reader to Appendix D.
@@
Our third local strategy is a greedy strategy, a sequential strategy that uses forward communication in order to optimally distinguish among the most likely hypothesis at every step. In this strategy each party $n\in\{1,\ldots,N\}$, starting with $n=1$, performs a measurement on the $n$-th subsystem of $\left|\psi_{k}\right\rangle\in\mathcal{H}^{\otimes N}$ corresponding to their location and attempts to identify the location of the error. As we show in Appendix D the optimal strategy for each party $n$ corresponds to a binary hypothesis testing scenario where the hypotheses are that the error is at location $n$, or at the most likely position among the remaining hypotheses given the prior information at party $n$’s disposal, by means of the prior distributions $p^{(n)}(k),\ k\in(1,\ldots N)$. These probabilities are updated sequentially, using Bayes’ rule, depending on the outcomes of all parties up to and including party $n-1$. Party $N$ then measures their subsystem using the optimal measurement derived from the prior distribution $p^{(N)}(k)$. The corresponding probability of success, which depends on the measurement record $\mathbf{m}\in\{0,1\}^{N-1}$, is $P^{(N)}_{S}(\mathbf{m})$ (see Eq. (24)) and the average probability of success is given by
@@
It is worth stressing that, aside from being easily implementable, there are other advantages that one needs to consider when choosing among local strategies. For example, the basic local strategy may identify the position of an error with high confidence in an online fashion, even in settings where the number of samples $N$, is not fixed in advance. The use of quantum sequential methodologies in settings without a finite horizon has been recently studied in Vargas et al. (2021); Li et al. (2022); Fanizza et al. (2023); Gasbarri et al. (2024).

### v3

@@ I Introduction
Recent advancements in quantum technologies, such as quantum computing devices [1, 2, 3, 4, 5, 6], quantum communication [7, 8, 9], and quantum sensors [10], lend credence to the notion that one day soon such devices will be readily available and, hopefully, part of an interconnected quantum network [11]. In turn, the existence of such quantum networks gives rise to new technical challenges, such as the correct identification of possible malfunctions. In a vast network of quantum devices—be they sources that produce quantum states, quantum channels that transmit information, or the vast array of gates in quantum computers—it is imperative that we are able to find efficient ways to identify faulty components.
@@
The identification of rare events that differ significantly from the majority of all other observations is known as anomaly detection and is a fundamental primitive in classical data analysis and signal processing [12] with a wide range of applications from the identification of denial of service attacks [13, 14, 15], to the identification of fraudulent financial activity [16] (for a survey of anomaly detection see [17]). Two important algorithms for anomaly detection, namely kernel principal component analysis and support vector machines, have been shown to be efficiently applicable in detecting anomalies in quantum data [18]. Here we consider a more direct application of anomaly detection where the anomaly is not restricted to classical data but in the most general quantum device. Specifically, given $N$ identical devices programmed to perform a particular task, with one of the devices developing a known malfunction, our goal is to optimally identify the malfunctioning device whilst only allowed to query the network once (see Fig. 1). This task that we term Error Position Identification, or EPI for short, is a key ingredient in pulse-position modulated quantum communication [19, 20], quantum illumination [21], quantum reading [22], and target detection [23].
@@
We note that instances of EPI have appeared elsewhere in the literature under the name of channel position finding [24, 25, 26, 27] where the task is to identify the position of a target bosonic channel, with a given reflectivity, among $m$ background bosonic channels of a different reflectivity. The case of identifying the position of a singular unitary channel was also considered recently in [28]. Here, we provide a more comprehensive analysis of EPI by considering more general devices—including states, unitary gates, and quantum channels—using a variety of different strategies employing probes in separable, entangled, as well as ancilla assisted strategies. Unlike [24, 25, 26, 27] we consider channels acting on two-dimensional systems, and consider both rank-one and rank-two Pauli noise channels, depolarizing, and amplitude damping channels. Moreover, our aim is to identify the position of the erroneous device by querying the network only once, as opposed to the more general multi-querying adaptive strategies considered in [24]. For the case of unitary channels we show that there exists instances for which the position of the unitary channel can be identified with certainty, a result that was missed in [28]. Note that EPI is fundamentally different from the anomaly detection scenario considered in [18] which focused in classification of states using quantum machine learning techniques.
@@
We first address the setting when the devices are quantum sources known to produce a specific pure state in $\mathcal{H}_{d}$ with the faulty source producing a different anomalous pure quantum state (the multi-anomaly case has recently been studied in [30]). We show that the maximum probability of success is achieved by a global measurement strategy given by the so-called square root measurement, which yields a success probability that converges to a non-zero constant as $N$ increases. We next show that when the devices are known to perform a specific unitary gate, with the anomalous device performing a different unitary, a probe state with entanglement among the $N$ input states—that can be effectively taken to be qubits—achieves the maximal probability of success without the need of using ancillary systems. In fact, we show that for certain non-trivial errors the optimal probability of success can even reach unity. Whilst the use of entangled probe states does yield an improvement over strategies employing separable probe states, this improvement diminishes with the size of the network.
@@
The task of successfully identifying the position of a faulty device is equivalent to optimally discriminating among $N$ quantum states, in the case of source EPI, or $N$ quantum channels. For the remainder of this work we will be concerned with devices that either prepare systems in a known pure state (sources), or perform a known quantum operation (channels). Without loss of generality a faultless quantum source prepares systems in the state $\left|0\right\rangle$ whereas the faulty source—equally likely to be located at any of the $N$ positions—prepares the $k^{\mathrm{th}}$ system in some state $\left|\phi\right\rangle$. Similarly, a faultless quantum operation performs the unitary operation $V$ whereas the anomalous operation performs a completely-positive, trace-preserving (CPTP) map, $\mathcal{F}^{(k)}$, on the $k^{\mathrm{th}}$ quantum system  [^1] In the language of [18] EPI is an instance of point anomaly and, moreover, can be cast in the framework of supervised learning as we assume that the description of both normal and anomalous devices is known. The probability of successfully identifying the position of the faulty device is given by
@@
where $q(l|k)$ is the conditional probability given by $\mathrm{Tr}\left[M_{l}\,\left|\phi_{k}\right\rangle\!\!\left\langle\phi_{k}\right|\right]$ in the case of sources, and $\mathrm{Tr}\left[M_{l}\,\mathcal{E}^{(k)}[\left|\psi\right\rangle\!\!\left\langle\psi\right|]\right]$ in the case of channels, with $\{M_{l}\geq 0\,|\sum_{l}M_{l}={\mbox{$1{\bf l}$}}\}:=\mathcal{M}$ forming a general positive operator-valued measure (POVM). Observe that for source EPI, the optimization over all measurement strategies of Eq. 1 is a semi-definite program (SDP). For channel EPI, however, Eq. 1 needs to be optimized both over the input state as well as the measurement. Luckily enough, the formalism of quantum testers [32] allows one to linearize the problem and formulate it as an SDP—as long as the optimization is carried over strategies that make use of idler or ancillary systems as shown in Fig. 1.
By and large analytical solutions to either state or channel discrimination problems are very difficult and known only for two [33, 34] or three [35] pure or mixed states or between states possessing a certain symmetry [36, 37, 38, 39, 40, 41, 42, 43, 44] (see also [45, 46]). For channel discrimination analytic results are known for distinguishing among a finite number of unitaries in a single or finite number of runs [47, 48, 49], or between two arbitrary CPTP maps [50, 51, 52]. We stress that here, unlike most channel discrimination instances to date, the identity of the channel (unitary or otherwise) is known, and what one is looking for is the position at which the channel is acting.
@@ III.1 Source EPI
Notice that the set of states also enjoys cyclic symmetry, which allows to work out the solution explicitly [37, 38, 39, 20, 30].
Let $\{\left|m_{j}\right\rangle\}_{j=1}^{N}$ be an orthonormal basis for the $N$-dimensional space spanned by $\{\left|\psi_{k}\right\rangle\}_{k=1}^{N}$. Then the probability of success is given by [42]
@@
where $S$ is the unique, self-adjoint square root of $G$. The corresponding measurement is simply given by $\{\left|m_{k}\right\rangle=\frac{(VS)^{-1}}{\sqrt{N}}\left|\psi_{k}\right\rangle\}$ (see [42]).
@@
where $b=\left\langle\psi_{k}\lvert\psi_{l}\right\rangle$ for $k\neq l$ and $\left|\bm{1}\right\rangle=\frac{1}{\sqrt{N}}(1,\ldots,1)^{T}$. The matrix $G$ and has two distinct eigenvalues, $\lambda_{1}=\frac{1+(N-1)b}{N}$, and $\lambda_{2}=\frac{1-b}{N}$ where the latter is $(N-1)$-fold degenerate. Note that $G$ is circulant [53] as is any function of $G$, in particular $S=\sqrt{G}$ whose diagonal entries are all equal to
@@
The condition that $S_{kk}=S_{ll},\,\forall k\neq l$ is necessary and sufficient to show that Eq. (3) is maximized by $B=S$ [20, 42], i.e., $P_{S}^{*}=NS_{kk}^{2}=(\mathrm{Tr}{S})^{2}/N$, which reads
@@
The measurement that achieves this optimal value is the so-called square-root measurement [37, 38, 39, 20]. Notice that $P_{S}^{*}=1$ if and only if $b=0$, i.e., $\left|\phi\right\rangle=\left|1\right\rangle$. For large strings of states, the success probability is finite and reads
@@ III.2 Unitary EPI
where $U=V^{\dagger}W$, and $\left|\psi\right\rangle\in\mathcal{H}_{p}\otimes\mathcal{H}_{a}$ is a probe state which in principle can contain ancillary systems. However, we note that the use of ancillas here is redundant since as for any probe-plus-ancilla state there exists an $N$ probe state that gives the same Gram matrix [54]. Observe that if $\left|\psi\right\rangle=\left|\gamma\right\rangle^{\otimes N}$ with $U\left|\gamma\right\rangle=\left|\phi\right\rangle$, we recover the source scenario discussed above with success probability given by Eq. (7). The question is whether choosing a more suitable initial state, perhaps involving entanglement, improves the probability of success.
Using the symmetry of the problem, for any given input state $\rho\in\mathcal{B}(\mathcal{H}^{\otimes N})$ and optimal POVM $\{M_{\sigma}\}$ there exists a permutationally invariant state $\tilde{\rho}\in\mathcal{B}(\mathcal{H}^{\otimes N})$ and permutationally covariant POVM $\{\tilde{M}_{\sigma}\}$ that achieves the same probability of success [54]. On the other hand, as the probability of success is a convex function the optimal input state can always be chosen to be pure. A seemingly natural, though unproven, assumption is to restrict the optimization over states to those that are both pure and permutationally invariant. For $N>2$ this assumption amounts to searching over pure states that belong to the totally symmetric subspace of $\mathcal{H}^{\otimes N}$ (see Eq. 11). We have numerically verified that this is indeed the case for small values of $N$.
As $U$ is known we may write any permutationally invariant probe state with respect to the eigenbasis of $U$, with all amplitudes taken to be real and positive without loss of generality. Furthermore, an optimal probe state only involves those eigenstates of $U$ whose eigenvalues have the largest distance in the complex plane [55, 56] as these yield the smallest overlap. We write these as $\left|0\right\rangle$ and $\left|1\right\rangle$, respectively. The action of the unitary can be taken to be $U\left|0\right\rangle=\left|0\right\rangle$ and $U\left|1\right\rangle=e^{i\phi}\left|1\right\rangle$ without loss of generality, where $|\phi|$ is the largest phase distance among the eigenvalues of the unitary. To keep track of the phase we denote the unitary by $U(\phi)$.
An orthonormal basis of these permutationally symmetric states is given by the well known Dicke states [57],
@@
We would like to point out that part of our results here are reproduced in a recent publication [28]. Notwithstanding, this work assumes symmetric probe states of qubits ab initio and fails to notice that this probability can be made equal to one for non-trivial values of the angle $\phi$.
@@ III.3 Channel EPI
For rank-three Pauli channels, any strategy involving $N$ probes prepared in a product state and projective measurements is sub-optimal. This is because one cannot perfectly discriminate between the Pauli matrices using a single qubit. Indeed, the best one can hope for using such a strategy is [54]
@@
We note that upper and lower bounds for the probability of success can be derived by suitably modifying the approach of [24] to the single shot scenario but that these bounds are not tight and are also not demonstrably known to be achievable.
@@ IV Local Measurement Strategies for source EPI
In this section, we investigate four local measurement strategies and gauge their performance in determining the position of a malfunctioning quantum source, i.e. detecting faulty prepared states. For ease of exposition, we will describe each of the local measurement strategies and state their performance only. For a more detailed treatment of each strategy, we direct the reader to the supplementary material [54].
@@
Our third local strategy is a greedy strategy, a sequential strategy that uses forward communication in order to optimally distinguish among the most likely hypothesis at every step. In this strategy each party $n\in\{1,\ldots,N\}$, starting with $n=1$, performs a measurement on the $n$-th subsystem of $\left|\psi_{k}\right\rangle\in\mathcal{H}^{\otimes N}$ corresponding to their location and attempts to identify the location of the error. As we show in [54] the optimal strategy for each party $n$ corresponds to a binary hypothesis testing scenario where the hypotheses are that the error is at location $n$, or at the most likely position among the remaining hypotheses given the prior information at party $n$’s disposal, by means of the prior distributions $p^{(n)}(k),\ k\in(1,\ldots N)$. These probabilities are updated sequentially, using Bayes’ rule, depending on the outcomes of all parties up to and including party $n-1$. Party $N$ then measures their subsystem using the optimal measurement derived from the prior distribution $p^{(N)}(k)$. The corresponding probability of success, which depends on the measurement record $\mathbf{m}\in\{0,1\}^{N-1}$, is $P^{(N)}_{S}(\mathbf{m})$ (see Eq. (24)) and the average probability of success is given by
@@
It is worth stressing that, aside from being easily implementable, there are other advantages that one needs to consider when choosing among local strategies. For example, the basic local strategy may identify the position of an error with high confidence in an online fashion, even in settings where the number of samples $N$, is not fixed in advance. The use of quantum sequential methodologies in settings without a finite horizon has been recently studied in [58, 59, 60, 61].